OpenStudy (anonymous):

HELPPPP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

PLZ

OpenStudy (anonymous):

HELP !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

wat do u think it is

OpenStudy (anonymous):

\[k^{-5}=\frac{1}{k^5}=\frac{1}{32}\] which tells you \[k^5=32\]

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

u need help with anyone of those

OpenStudy (anonymous):

yo u need help with anymore of those questions or not

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

THERE LOL

OpenStudy (thewalkingdeadrick):

what grade are you in

OpenStudy (anonymous):

8

OpenStudy (anonymous):

-3n<81 -3n/-3<81/3 n>-27

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

np

OpenStudy (madhu.mukherjee.946):

@Peaches__15 great job

OpenStudy (anonymous):

@madhu.mukherjee.946 thx

OpenStudy (anonymous):

@ajkishon09 lol ur cute

OpenStudy (anonymous):

@ajkishon09 do u have more question just post dem and ill answer dem k

OpenStudy (anonymous):

@ajkishon09 u know u can close dis question